A sequence {a} is defined as follows: 20 = 2.0, = 1, and for n > 2,0 = 3.27-1-nan-2 +1. What is az? 2 -1 -2 1
Answer #1

an = 3 an-1- nan-at) Az = 32 - 3. A, +) a = 3 ay-200+) - 3.1-2.2+1 3-4+1 = 0. -3.0-3.14 - - 3 t1 = -2 o